The probability of a successful missile launch is 0.83. Test launches are conducted until three successful launches are achieved. What is the probability that

(c) At least 6 launches will be required?

(d) Will your answer change if we require three successful launches to be consecutive? No need to do the calculations again, but make sure to explain your answer.
Answer #1

c)

P( At least 6 launches will be required )=P(in first 5 launches at most 2 successful missile launch )

=P(X=0)+P(X=1)+P(X=2)=5C0(0.83)0(0.17)5+5C1(0.83)1(0.17)4+5C2(0.83)2(0.17)3=0.0375

d)

as having three successful launches to be consecutive will add another condition on above event which will reduce the number of ways above event can happen, and therefore the answer will change cause the number of ways the combined event can happen does change.,